OpenStudy (anonymous):

Probability question P(n,5)=12x P(n,3))

OpenStudy (anonymous):

P(n,5)=12 *P(n,3) \[n \left( n-1 \right)\left( n-2 \right)\left( n-3 \right)\left( n-4 \right)=12 n \left( n-1 \right)\left( n-2 \right)\] \[\left( n-3 \right)\left( n-4 \right)=12,n ^{2}-4n-3n+12=12,n ^{2}-7n=0,n \left(n-7 \right)=0\] rejecting n=0, hence n=7